# FraudTransformer: Time-Aware GPT for Transaction Fraud Detection

### v1

@@ 1. Introduction
Based on these findings we propose an algorithm which we call FraudTransformer, combining a vanilla GPT architecture with relative sinusoidal embeddings.
@@ 5. Encoding process
We use a learnable lookup table $W_{\text{tok}}\!\in\!\mathbb{R}^{V\times d_{\text{model}}}$ to map token id $x_{t}\!\in\!\{0,\ldots,V{-}1\}$ to a vector:
@@
This contributes $V\!\cdot\!d_{\text{model}}$ parameters and is trained end-to-end. We tie weights with the LM head (the output projection reuses $W_{\text{tok}}^{\top}$), so the encoder does not introduce an additional vocabulary-scale matrix.
@@
$e^{\text{col}}_{t}\;=\;W_{\text{col}}[c_{t}],\qquad W_{\text{col}}\!\in\!\mathbb{R}^{C\times d_{\text{model}}},$
where $c_{t}\!\in\!\{0,\ldots,C{-}1\}$ indexes the token’s source column and $C$ is the number of columns. This adds $C\!\cdot\!d_{\text{model}}$ parameters and is learned jointly with the model. Intuitively, $e^{\text{col}}_{t}$ provides a field-specific signal that helps the attention layers learn different patterns for different columns of features, even when the raw tokens overlap.
@@ 6. Algorithm and Model
We implement a vanilla GPT-style architecture (radford2018improving, ) with $8$ heads, $6$ layers, context length ($L_{\max}$) $1024$, embedding size ($d_{\text{model}}$) of $512$, resulting in a model with around $20M$ parameters. A summary of model configuration is provided in Table 2.
@@
Embedding size ($d_{\text{model}}$) $512$
@@
We let $d_{\text{model}}$ denote the GPT model width (the residual dimension), and $L_{\max}$ denote the maximum sequence length (context length) of the absolute position table for the positional encoder. For vectors $a,b\in\mathbb{R}^{n}$, the Hadamard product is $a\odot b=(a_{1}b_{1},\dots,a_{n}b_{n})$.
Let $t$ index tokens and $l$ index layers. Let $h_{t}^{(l)}\in\mathbb{R}^{d_{\text{model}}}$ be the residual (pre-attention) hidden state in layer $l$ for token $t$. Define a discrete time-id $\mathrm{id}_{t}\in\{0,1,2,\ldots\}$ from wall-clock time or offsets:
@@
where $t_{\text{event}}{=t_{\text{event}}(t)}$ is the time of event to which the token $t$ belongs. Let $p_{t}\in\{0,1,\dots\}$ be the position index of token $t$; often, this is different from $t_{\text{event}}(t)$.
@@
    Time encoding (additive features). For a given base frequency $f_{\rm base}$, which is a hyperparameter, we build $e_{t}\in\mathbb{R}^{d_{\text{model}}}$ from $\mathrm{id}_{t}$ using one of the two options below.
    • Sinusoidal encoding (GPT-style Fourier features): for even $d_{\text{model}}$, let $\theta_{i}=f_{\text{base}}^{-2i/d_{\text{model}}}$ with base frequency $f_{\text{base}}$ and $i=0,\dots,\tfrac{d_{\text{model}}}{2}-1$, and set
      $e_{t}^{\sin}=\big[\,\sin(\theta\odot\mathrm{id}_{t})\ ;\ \cos(\theta\odot\mathrm{id}_{t})\,\big]\in\mathbb{R}^{d_{\text{model}}}.$
    • Rotary-derived (used additively): Assume $d_{\text{model}}$ is even. Define the inverse-frequency values with base frequency $f_{\text{base}}$:
      $\omega_{i}\;=\;f_{\text{base}}^{-\,\frac{2i}{d_{\text{model}}}},\qquad i=0,\dots,\frac{d_{\text{model}}}{2}-1.$
@@
      Concatenating these pairs over $i=0,\dots,\frac{d_{\text{model}}}{2}-1$ in alternating order (even/odd slots) yields $e_{t}\in\mathbb{R}^{d_{\text{model}}}$.
@@
    (1) $\begin{split}\mu(e_{t})&=\frac{1}{d_{\text{model}}}\sum_{i=1}^{d_{\text{model}}}e_{t,i},\qquad\\ \sigma^{2}(e_{t})&=\frac{1}{d_{\text{model}}}\sum_{i=1}^{d_{\text{model}}}\bigl(e_{t,i}-\mu(e_{t})\bigr)^{2}.\end{split}$
@@
    Learned Positional Encoders. We use GPT-2–style learnable positional embeddings (hf-gpt2-docs-v4-38-0, ). We maintain a parameter matrix $\mathrm{EmbPos}\in\mathbb{R}^{L_{\max}\times d_{\text{model}}}$ and perform a row lookup by position index $p_{t}\in\{0,\dots,L_{\max}-1\}$:
@@
    where $e^{\text{tok}}_{t},e^{\text{col}}_{t}\in\mathbb{R}^{d_{\text{model}}}$ are token and column embeddings, respectively and $p_{t}^{\mathrm{learn}},\hat{e}_{t}\in\mathbb{R}^{d_{\text{model}}}$. The Transformer blocks then proceed unchanged with $\bar{h}_{t}^{(1)}$ as input embedding.
@@ 7.4. Fraud Types
7.5. Ablation Study
ID Time encoder Time scheme Positional encoder
SA Sinusoidal Absolute
SAP Sinusoidal Absolute
SR Sinusoidal Relative
SRP Sinusoidal Relative
RA Rotary Absolute
RAP Rotary Absolute
$R$ Rotary Relative
RRP Rotary Relative
7.6. Results
Table 5. Baseline models and Transformer variants evaluated on the HSBC Payment Fraud Group’s test set (mean $\pm$ standard deviation over three runs; higher is better). TE=time encoder, PE=positional encoder. Time–encoder codes follow the scheme introduced in Table 4: S=sinusoidal, R=rotary; A=absolute timestamps, R=relative.
Model PRAUC $\uparrow$ AUROC $\uparrow$
Feature-based baselines
 Logistic Regression $0.92491\pm 0.00000$ $0.93784\pm 0.00000$
 XGBoost $0.93447\pm 0.00000$ $0.95894\pm 0.00000$
 LightGBM $0.94536\pm 0.00000$ $0.96212\pm 0.00000$
Transformer – no positional encoder
 Vanilla (no TE, no PE) $0.92038\pm 0.00019$ $0.93602\pm 0.00067$
RA (rotary, Abs) $0.91278\pm 0.00010$ $0.93003\pm 0.00021$
RR (rotary, Rel) $0.92145\pm 0.00049$ $0.93989\pm 0.00072$
SA (sinusoidal, Abs) $0.91539\pm 0.00043$ $0.93034\pm 0.00079$
SR (sinusoidal, Rel) $0.92477\pm 0.00012$ $0.94212\pm 0.00067$
Transformer – with positional encoder
NTE (no TE, with PE) $0.94689\pm 0.00109$ $0.95664\pm 0.00064$
RAP (rotary, Abs) $0.94436\pm 0.00080$ $0.95539\pm 0.00022$
RRP (rotary, Rel) $0.95610\pm 0.00118$ $0.96497\pm 0.00041$
SAP (sinusoidal, Abs) $0.94713\pm 0.00042$ $0.95639\pm 0.00093$
SRP (sinusoidal, Rel) $\boldsymbol{0.95816\pm 0.00092}$ $\boldsymbol{0.96723\pm 0.00011}$
SRP – without LayerNorm (WOL)
 SRP/WOL $0.95304\pm 0.00102$ $0.96301\pm 0.00091$
Table 5 contrasts the proposed GPT with Time encoder and Positional encoder model with the three classical feature-based learners from Section 7.2 and with different reduced GPT variants. All runs used fixed train/validation/test splits and deterministic hyperparameters  [^1] We select no row/feature subsampling: LightGBM bagging_fraction=1.0, feature_fraction=1.0; XGBoost subsample=1.0, colsample_bytree=1.0.. Consequently, training and evaluation are deterministic, yielding identical results across repeated runs. Features are from a single time snapshot for these comparisons. Accordingly, the standard deviation of the results for the three baselines are zero. The results are shown in Table 5.
@@ 7.7. Discussion
Our strongest configuration is sinusoidal time embeddings with learned positional embeddings, event-level Relative, and LayerNorm. In the following, we provide a more detailed discussion based on our observations in experiments.
@@ 8. Conclusion
In practice, FraudTransformer remains computationally lightweight, requiring only a single training epoch for hyperparameter sweeps, and delivers stable improvements that translate into fewer missed frauds and lower false-alarm rates.
DISLCAIMER
This paper was prepared for information purposes, and is not a product of HSBC or its affiliates. Neither HSBC nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper but not limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. Copyright HSBC Group 2023. The content of this manuscript was reviewed by HSBC’s publication clearance team to ensure no disclosure of sensitive data.

### v2

@@ 1. Introduction
Based on these findings we propose an algorithm which we call FraudTransformer, combining a vanilla GPT architecture with time and positional encoders.
@@ 5. Encoding process
In the following, we let $d_{\text{emb}}$ denote the embedding size of GPT model.
@@
We use a learnable lookup table $W_{\text{tok}}\!\in\!\mathbb{R}^{V\times d_{\text{emb}}}$ to map token id $x_{t}\!\in\!\{0,\ldots,V{-}1\}$ to a vector:
@@
This contributes $V\!\cdot\!d_{\text{emb}}$ parameters and is trained end-to-end. We tie weights with the LM head (the output projection reuses $W_{\text{tok}}^{\top}$), so the encoder does not introduce an additional vocabulary-scale matrix.
@@
$e^{\text{col}}_{t}\;=\;W_{\text{col}}[c_{t}],\qquad W_{\text{col}}\!\in\!\mathbb{R}^{C\times d_{\text{emb}}},$
where $c_{t}\!\in\!\{0,\ldots,C{-}1\}$ indexes the token’s source column and $C$ is the number of columns. This adds $C\!\cdot\!d_{\text{emb}}$ parameters and is learned jointly with the model. Intuitively, $e^{\text{col}}_{t}$ provides a field-specific signal that helps the attention layers learn different patterns for different columns of features, even when the raw tokens overlap.
@@ 6. Algorithm and Model
We let $L_{\max}$ denote the maximum sequence length (context length). We implement a vanilla GPT-style architecture (radford2018improving, ) with $8$ heads, $6$ layers, context length ($L_{\max}$) of $1024$, and embedding size ($d_{\text{emb}}$) of $512$, resulting in a model with around $20M$ parameters. A summary of the model configuration is provided in Table 2.
@@
Embedding size ($d_{\text{emb}}$) $512$
@@
Let $t$ index tokens and $l$ index layers. Let $h_{t}^{(l)}\in\mathbb{R}^{d_{\text{emb}}}$ be the residual (pre-attention) hidden state in layer $l$ for token $t$. Define a discrete time-id $\mathrm{id}_{t}\in\{0,1,2,\ldots\}$ from wall-clock time or offsets:
@@
where $t_{\text{event}}{=t_{\text{event}}(t)}$ is the time of event to which the token $t$ belongs. Let $p_{t}\in\{0,1,\dots\}$ be the position index of token $t$; often, this is different from $t_{\text{event}}(t)$. For vectors $a,b\in\mathbb{R}^{n}$, the Hadamard product is $a\odot b=(a_{1}b_{1},\dots,a_{n}b_{n})$.
@@
    Time encoding (additive features). For a given base frequency $f_{\rm base}$, which is a hyperparameter, we build $e_{t}\in\mathbb{R}^{d_{\text{emb}}}$ from $\mathrm{id}_{t}$ using one of the two options below.
    • Sinusoidal encoding (GPT-style Fourier features): for even $d_{\text{emb}}$, let $\theta_{i}=f_{\text{base}}^{-2i/d_{\text{emb}}}$ with base frequency $f_{\text{base}}$ and $i=0,\dots,\tfrac{d_{\text{emb}}}{2}-1$, and set
      $e_{t}^{\sin}=\big[\,\sin(\theta\odot\mathrm{id}_{t})\ ;\ \cos(\theta\odot\mathrm{id}_{t})\,\big]\in\mathbb{R}^{d_{\text{emb}}}.$
    • Rotary-derived (used additively): Assume $d_{\text{emb}}$ is even. Define the inverse-frequency values with base frequency $f_{\text{base}}$:
      $\omega_{i}\;=\;f_{\text{base}}^{-\,\frac{2i}{d_{\text{emb}}}},\qquad i=0,\dots,\frac{d_{\text{emb}}}{2}-1.$
@@
      Concatenating these pairs over $i=0,\dots,\frac{d_{\text{emb}}}{2}-1$ in alternating order (even/odd slots) yields $e_{t}\in\mathbb{R}^{d_{\text{emb}}}$.
@@
    (1) $\begin{split}\mu(e_{t})&=\frac{1}{d_{\text{emb}}}\sum_{i=1}^{d_{\text{emb}}}e_{t,i},\qquad\\ \sigma^{2}(e_{t})&=\frac{1}{d_{\text{emb}}}\sum_{i=1}^{d_{\text{emb}}}\bigl(e_{t,i}-\mu(e_{t})\bigr)^{2}.\end{split}$
@@
    Learned Positional Encoders. We use GPT-2–style learnable positional embeddings (hf-gpt2-docs-v4-38-0, ). We maintain a parameter matrix $\mathrm{EmbPos}\in\mathbb{R}^{L_{\max}\times d_{\text{emb}}}$ and perform a row lookup by position index $p_{t}\in\{0,\dots,L_{\max}-1\}$:
@@
    where $e^{\text{tok}}_{t},e^{\text{col}}_{t}\in\mathbb{R}^{d_{\text{emb}}}$ are token and column embeddings, respectively and $p_{t}^{\mathrm{learn}},\hat{e}_{t}\in\mathbb{R}^{d_{\text{emb}}}$. The Transformer blocks then proceed unchanged with $\bar{h}_{t}^{(1)}$ as input embedding.
ID Time encoder Time scheme Positional encoder
SA Sinusoidal Absolute
SAP Sinusoidal Absolute
SR Sinusoidal Relative
SRP Sinusoidal Relative
RA Rotary Absolute
RAP Rotary Absolute
$R$ Rotary Relative
RRP Rotary Relative
[…]
Model PRAUC $\uparrow$ AUROC $\uparrow$
[…]
Feature-based baselines
 Logistic Regression $0.92491\pm 0.00000$ $93784\pm 0.00000 +0.0$0453+0.00453 +0.00182+0.00182
 XGBoost $0.93447\pm 0.00000$ $95894\pm 0.00000 +0.0$1409+0.01409 +0.02292+0.02292
 LightGBM $0.94536\pm 0.00000$ $96212\pm 0.00000 +0.0$2498+0.02498 +0.02610+0.02610
Transformer – no positional encoder
 Vanilla (baseline)  (no TE, no PE) $0.92038\pm 0.00019$ $93602\pm 0.00067 +0.0$0000+0.00000 +0.00000+0.00000
RA (rotary, Abs) $0.91278\pm 0.00010$ $93003\pm 0.00021 0.0$0760-0.00760 0.00599-0.00599
RR (rotary, Rel) $0.92145\pm 0.00049$ $93989\pm 0.00072 +0.0$0107+0.00107 +0.00387+0.00387
SA (sinusoidal, Abs) $0.91539\pm 0.00043$ $93034\pm 0.00079 0.0$0499-0.00499 0.00568-0.00568
SR (sinusoidal, Rel) $0.92477\pm 0.00012$ $94212\pm 0.00067 +0.0$0439+0.00439 +0.00610+0.00610
Transformer – with positional encoder
NTE (no TE, with PE) $0.94689\pm 0.00109$ $95664\pm 0.00064 +0.0$2651+0.02651 +0.02062+0.02062
RAP (rotary, Abs) $0.94436\pm 0.00080$ $95539\pm 0.00022 +0.0$2398+0.02398 +0.01937+0.01937
RRP (rotary, Rel) $0.95610\pm 0.00118$ $96497\pm 0.00041 +0.0$3572+0.03572 +0.02895+0.02895
SAP (sinusoidal, Abs) $0.94713\pm 0.00042$ $95639\pm 0.00093 +0.0$2675+0.02675 +0.02037+0.02037
SRP (sinusoidal, Rel) $\boldsymbol{0.95816\pm 0.00092}$ $\boldsymbol{0.96723\pm 0.00011}$ $\boldsymbol{+0.03778}$ $\boldsymbol{+0.03121}$
SRP – without LayerNorm (WOL)
 SRP/WOL $0.95304\pm 0.00102$ $96301\pm 0.00091 +0.0$3266+0.03266 +0.02699+0.02699
@@ 7.6. Results
Table 4 contrasts the proposed GPT with Time encoder and Positional encoder model with different positional encoders with the three classical feature-based learners from Section 7.2 and with different reduced GPT variants. All runs used fixed train/validation/test splits and deterministic hyperparameters  [^1] We select no row/feature subsampling: LightGBM bagging_fraction=1.0, feature_fraction=1.0; XGBoost subsample=1.0, colsample_bytree=1.0.. Consequently, training and evaluation are deterministic, yielding identical results across repeated runs. Features are from a single time snapshot for these comparisons. Accordingly, the standard deviation of the results for the three baselines are zero. The results are shown in Table 4.
@@ 7.7. Discussion
Our strongest configuration is sinusoidal time embeddings with learned positional embeddings, event-level Relative, and LayerNorm, appearing as SRP in Table 4. In the following, we provide a more detailed discussion based on our observations in experiments.
PRAUC Values: The PRAUC baseline is the proportion of positive (that is here, fraudulent) entries. Hence with a $43\%$ fraud rate in our evaluation set, the random baseline is $0.43$, which may make the PRAUC values appear to be high. Therefore, instead of reporting absolute PRAUC, we emphasise relative differences between models on the same test set. For this purpose the $\Delta$ of PRAUC with respect to baseline, Vanilla GPT, is reported in Table 4.
@@ 8. Conclusion and Future Works
In practice, FraudTransformer remains computationally lightweight, requiring only a single training epoch for hyperparameter sweeps. Moreover, FraudTransformer delivers stable improvements that translate into fewer missed frauds and lower false-alarm rates.
Some improvements are possible. This work trains the model end-to-end with a binary cross-entropy objective for sequence-level fraud classification. In future work, we suggest to pre-train the GPT backbone using a next-token prediction objective (wang2024emu3, ), with the expectation that the token positions will be learnt more accurately. While our current task is binary fraud vs. non-fraud, we could extend the model to multi-class prediction over fraud subtypes in future works. Finally, our approach can be extended to new GPT architectures which support larger context lengths.
DISCLAIMER
This paper was prepared for information purposes, and is not a product of HSBC or its affiliates. Neither HSBC nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper but not limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. Copyright HSBC Group 2025. The content of this manuscript was reviewed by HSBC’s publication clearance team to ensure no disclosure of sensitive data.